# The FFT Strikes Back: An Efficient Alternative to Self-Attention

### v4

@@ Abstract
Conventional self-attention mechanisms incur quadratic complexity, limiting their scalability on long sequences. We introduce FFTNet, an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time. By transforming inputs into the frequency domain, FFTNet exploits the orthogonality and energy preservation guaranteed by Parseval’s theorem to capture long-range dependencies efficiently. Our main theoretical contributions are 1) an adaptive spectral filter, 2) combining local windowing with a global FFT branch, and 3) rich nonlinearity introduction in both the frequency and token domains. Experiments on the Long Range Arena and ImageNet benchmarks validate our theoretical insights and demonstrate superior performance over fixed Fourier and standard attention models.
@@ 1 Introduction
Conventional self-attention mechanisms capture global interactions through explicit pairwise computations, which results in a quadratic computational complexity that can be prohibitive for long sequences. In contrast, our work introduces an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to perform global token mixing with a mathematically elegant and scalable approach.
Our method begins by transforming the input sequence into the frequency domain, where orthogonal frequency components naturally encode long-range dependencies. This not only reduces the computational complexity to $\mathcal{O}(n\log n)$ but also preserves the energy of the original signal, as ensured by Parseval’s theorem. Such a transformation facilitates efficient global interactions without the need for exhaustive pairwise comparisons.
A central innovation of our framework is the integration of a learnable spectral filter. This adaptive component modulates the Fourier coefficients based on a global context vector, enabling the model to dynamically emphasize salient frequency bands that are critical for capturing complex patterns. Furthermore, the application of nonlinear activations to both the real and imaginary parts of the filtered signal enhances the model’s expressivity, allowing it to represent higher-order interactions that go beyond the scope of linear transformations.
In essence, our adaptive spectral filtering framework combines the computational efficiency of FFT-based transformations with adaptive, context-sensitive filtering and nonlinear processing. This synthesis offers a rigorous and expressive alternative to traditional self-attention, providing a robust solution for modeling long-range dependencies in sequence data.
@@ 3.1 Motivation
Why Local Windowing?
While the global DFT branch is excellent for capturing large-scale dependencies, it can blur fine-grained or short-range signals. Local windowing (STFT) is optional, but it can help preserve the high-resolution structure within shorter segments. When used, merging these local and global representations yields a model capable of both wide-ranging context and precise local detail. If purely global context suffices, one may omit local windowing altogether.
@@ 3.2 Method Description
1. Global Fourier Transform.
@@
2. Local Windowing (STFT). (Optional)
In parallel, and optionally, perform a short-time Fourier transform (STFT) for local context:
@@
    Partition the input $\mathbf{X}$ into overlapping windows of length $w$. Let each window be denoted by $\mathbf{X}^{(m)}\in\mathbb{R}^{w\times d}$, where $m$ indexes the window.
@@
These local spectra emphasize short-range patterns that might be blurred in a purely global transform. If the application does not require high-resolution local detail, this step can be omitted.
3. Merging Global and Local Representations.
To maximize information flow from both branches, we propose an isometric fusion approach in the frequency domain. First, concatenate the global and local spectra along their feature dimension: [^1] If multiple local windows exist, they may be pooled or aggregated (e.g. average or max) before concatenation, or one can apply a deeper fusion network to combine them.
@@
where $\mathbf{F}_{\mathrm{local}}$ can be a representative fusion of all local windows. [^2] If retaining every window separately, this dimension becomes $n\times(d+\sum_{m}d_{\mathrm{local}}^{(m)})$.
@@
4. Adaptive Spectral Filtering.
@@
Here, $\mathbf{F}$ includes both local and global frequency information from the isometric fusion step.
5. Nonlinear Activation (modReLU) and Inverse FFT.
@@
Finally, we map back to the token domain via the inverse FFT:
$\mathbf{Y}\;=\;\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d},$
yielding a globally mixed representation that also incorporates local spectral features (if local windowing was used).
@@ 3.3 Efficient Mixing
Using the FFT to encode global interactions has several advantages. It performs global mixing in $\mathcal{O}(n\log n)$ time, contrasting with the $\mathcal{O}(n^{2})$ cost of self-attention. Meanwhile, local STFT windows capture short-range structure that may be lost in a purely global transform. By isometrically merging local and global frequency components, we fuse high-resolution local context with a global view, all in roughly $\mathcal{O}(n\log n)$ time.
@@ 3.6 Proof of Computational Complexity
Local Windowing (STFT). (Optional)
If each window has length $w$ and we assume an overlap of $\alpha w$ tokens, the number of windows is approximately $\frac{n}{(1-\alpha)w}$. Each windowed FFT thus costs $\mathcal{O}(w\log w)$ for each of the $d$ channels. The total cost is:
$\mathcal{O}\!\Bigl{(}\tfrac{n}{(1-\alpha)w}\;d\;w\log w\Bigr{)},$
which is typically smaller or comparable to $\mathcal{O}(d\,n\log n)$ for practical choices of $w$. This step is optional and can be skipped if local structure is not critical for the task at hand.
@@
Figure 1: Overall Process of FFTNet with Local Windowing (optional) and Global FFT, followed by isometric fusion in the frequency domain.
@@ 3.7 Comparison with Convolution-Based Methods
Although a sufficiently large convolution kernel can span the entire sequence, it generally lacks the adaptability of self-attention unless the kernel itself is dynamically generated. Our approach bridges this gap by using a data-dependent, learnable filter in the frequency domain, guided by global context and applied through modReLU. Because frequency-domain multiplication corresponds to convolution in the token domain, the model can realize global, context-aware mixing more flexibly than standard convolutional methods. Local windowing further refines the representation without resorting to large kernel sizes.
@@ 3.8 Comparison with FNet
FNet Lee-Thorp et al. (2022) similarly uses the DFT for token mixing but employs no learnable parameters in the mixing step. In contrast, our method includes a context-dependent, adaptive filter and a nonlinear activation that both enhance expressivity and allow emphasis on specific frequency components. Further, we incorporate a local STFT pathway to capture fine-grained interactions. Despite these adaptive steps, the overall complexity remains $\mathcal{O}(n\log n)$, dominated by the FFT itself.
@@ 3.8.1 Mitigating Complexity in the Adaptive MLP
For very long sequences, generating frequency-bin parameters can still be demanding. Possible solutions include (1) subsampling or pooling the frequency representation before the MLP, (2) sharing parameters among adjacent frequency bins, and (3) applying low-rank factorization to reduce dimensionality. These techniques help maintain a reasonable balance between the expressive power of the adaptive filter and computational cost.
Conclusion. Adaptive spectral filtering integrates frequency-domain global mixing, data-driven filtering, and nonlinear activation into a unified framework. By adding an optional local windowing path (STFT) and merging it with the global FFT via an isometric fusion, the model captures both short-range and long-range interactions with $\mathcal{O}(n\log n)$ efficiency. Parseval’s theorem and the orthogonal properties of the DFT ensure stable transformations, enabling many benefits of self-attention at a reduced computational cost. Full implementation details are available at: https://github.com/jacobfa/fft.
@@ 4 Experiments
We present a comprehensive evaluation of our proposed approach, FFTNet, comparing it against both FNet and traditional Transformers based on self-attention. Our experiments cover the Long Range Arena (LRA) benchmark Tay et al. (2021) and ImageNet classification. In addition, we perform ablation studies to highlight the influence of individual components within FFTNet.
@@ 4.2 Long Range Arena (LRA) Benchmark
We evaluate on six LRA tasks: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X. Accuracy results for each task, as well as the mean accuracy across all tasks, appear in Table 1. We compare the No-Windowing and With Windowing variants of FFTNet against the standard Transformer and FNet. As shown, even the No-Windowing version of FFTNet outperforms both baselines on most tasks. Adding local windowing (STFT) further improves the accuracy slightly.
Model ListOps Text Retrieval Image Pathfinder Path-X Avg. Transformer 36.06 61.54 59.67 41.51 80.38 OOM 55.83 FNet 35.33 65.11 59.61 38.67 77.80 FAIL 55.32 FFTNet (No-Windowing) 37.65 66.01 60.21 42.02 80.71 83.25 58.31 FFTNet (With Windowing) 38.02 66.25 60.64 42.45 80.99 83.64 58.83
Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark tasks. “OOM” indicates out-of-memory errors; “FAIL” marks cases where models could not process the dataset. FFTNet (With Windowing) achieves the highest average accuracy.
@@ 4.3 ImageNet Classification
We further assess our FFTNetViT variants on ImageNet, comparing them against standard ViT baselines. Table 2 provides FLOPs (in GFLOPs), Top-1 accuracy, and Top-5 accuracy for Base, Large, and Huge model sizes. Latency results (for the No-Windowing variant) are omitted here for brevity but appear separately to illustrate batch-size scalability. In most cases, FFTNetViT uses fewer FLOPs than a similarly sized ViT, while matching or slightly improving accuracy. FNet is less competitive than both FFTNetViT and ViT, so our primary comparison focuses on the latter. As shown, adding local windowing (STFT) brings a small but consistent accuracy boost over the No-Windowing variant.
FFTNetViT (No-Windowing) FFTNetViT (With Windowing) ViT Variant FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) FLOPs Top-1 (%) Top-5 (%) Base 22.64 79.6 $\uparrow 0.2\%$ 94.9 $\uparrow 0.1\%$ 22.64 79.8 $\uparrow 0.4\%$ 95.0 $\uparrow 0.2\%$ 36.65 79.4 94.8 Large 79.92 82.1 $\uparrow 0.3\%$ 96.2 $\uparrow 0.2\%$ 79.92 82.3 $\uparrow 0.5\%$ 96.3 $\uparrow 0.3\%$ 127.18 81.8 96.0 Huge 166.14 83.2 $\uparrow 0.3\%$ 96.8 $\uparrow 0.2\%$ 166.14 83.4 $\uparrow 0.5\%$ 96.9 $\uparrow 0.3\%$ 261.39 82.9 96.6
Table 2: Comparison on ImageNet for Base, Large, and Huge variants. FLOPs are given in GFLOPs. “No-Windowing” vs. “With Windowing” are two configurations of FFTNetViT. The green arrows indicate improvements over the corresponding ViT.
@@
Throughput and Speedup Analysis (No-Windowing).
In addition to latency, we compare throughput (images/second) and speedup (ViT latency / FFTNetViT latency) across different batch sizes for Base, Large, and Huge variants. Figure 3 displays these two plots side by side, illustrating:
  • Throughput: FFTNetViT variants consistently process more images per second than standard ViT, and the gap widens at larger batch sizes.
  • Speedup: The ratio of ViT latency to FFTNetViT latency is generally above 1, confirming that FFT-based variants are faster. This advantage becomes more pronounced as the batch size increases.
@@ 4.4 Ablation Studies
Ablation studies on ImageNet (Base variant, No-Windowing) were conducted to isolate key FFTNet components. Four configurations are examined. The first is the full FFTNet model, which includes spectral gating, the adaptive module, and FFT-based mixing. The second removes spectral gating. The third excludes the adaptive filtering mechanism. The fourth replaces FFT-based mixing with a standard convolution. Table 3 shows the Top-1 accuracy of each variant. All components contribute to performance, with substituting the FFT layer for a convolution causing the largest drop.
Variant Top-1 Acc (%) Observation FFTNet (full) 79.6 Contains spectral gating, adaptive module, and FFT-based filtering. Without spectral gating 78.3 Omitting gating causes a moderate performance decline. Without adaptive module 77.8 Disabling the adaptive module degrades accuracy further. Replace FFT with convolution 77.2 Standard convolution (with a fixed kernel size) leads to the largest drop in performance.
Table 3: Ablation experiments on ImageNet (Base variant, No-Windowing). Each row omits or replaces one key FFTNet component from the full setup.
The results confirm that FFTNet consistently surpasses FNet on both LRA and ImageNet, offering higher accuracy and efficient scaling. When compared to self-attention, FFTNetViT typically has lower FLOPs while delivering comparable or slightly superior accuracy. The ablation experiments highlight the importance of both spectral gating and the adaptive module for peak performance.

### v5

# The FFT Strikes Again: An Efficient Alternative to Self-Attention

@@ Abstract
Conventional self-attention mechanisms exhibit quadratic complexity in sequence length, making them challenging to scale for long inputs. We present FFTNet, an adaptive spectral filtering framework that uses the Fast Fourier Transform (FFT) to achieve global token mixing in $\mathcal{O}(n\log n)$ time. By mapping inputs into the frequency domain, FFTNet exploits orthogonality and energy preservation—guaranteed by Parseval’s theorem—to efficiently model long-range dependencies. Our main theoretical contributions include 1) An adaptive spectral filter that highlights salient frequency components, 2) A hybrid scheme combining local windowing with a global FFT branch, 3) Nonlinear feature transformations applied in both the frequency and token domains. Experiments on Long Range Arena and ImageNet validate our theoretical insights and demonstrate superior performance over fixed Fourier-based and standard attention models.
@@ 1 Introduction
Conventional self-attention mechanisms capture global interactions through explicit token-pair computations, resulting in quadratic complexity $\mathcal{O}(n^{2})$. This high cost becomes prohibitive for tasks involving lengthy sequences, such as long-context language modeling or large-scale image processing.
In this paper, we introduce an adaptive spectral filtering framework that leverages the Fast Fourier Transform (FFT) to perform global token mixing with $\mathcal{O}(n\log n)$ efficiency. By converting the input sequence into the frequency domain, we harness orthogonal frequency components to naturally capture long-range dependencies. This frequency-domain representation not only reduces computational overhead but also preserves the original signal’s energy due to Parseval’s theorem.
A distinguishing feature of our approach is a learnable spectral filter that modulates the Fourier coefficients based on a global context vector. This adaptive mechanism selects which frequency bands are most relevant, allowing the model to dynamically focus on critical components of the input. Additionally, introducing nonlinear activations to both the real and imaginary parts of the signal extends the representational capacity beyond linear transformations, enabling the model to capture more complex interactions.
Overall, our adaptive spectral filtering framework unifies the efficiency of FFT-based methods with a powerful, input-dependent spectral filter and nonlinear feature modulation. This combination offers a computationally efficient and expressive alternative to traditional self-attention, promising robust handling of long-range dependencies in various sequence modeling tasks.
@@ 2.2 Fourier-Based Mixing
In another line of work, Xu et al. Xu et al. (2020) propose learning-based frequency selection methods for large-scale vision tasks, showing that CNNs can be trained directly on partial frequency components while retaining high accuracy. Their approach offers a path to reducing data transmission overhead and computational costs without sacrificing performance.
@@ 3.1 Motivation
Why Local Windowing (or Wavelets)?
While the global DFT branch is excellent for capturing large-scale dependencies, it can blur fine-grained or short-range signals. Local windowing (STFT) is optional, but it can help preserve the high-resolution structure within shorter segments. In fact, wavelet transforms can also be used here instead of or alongside STFT, offering a multi-resolution perspective that naturally adapts to local signal changes. When used, merging these local (STFT or wavelets) and global representations yields a model capable of both wide-ranging context and precise local detail. If purely global context suffices, one may omit local windowing or wavelet transforms altogether.
@@ 3.2 Method Description
1
Global Fourier Transform.
@@
2
Local Windowing (STFT) or Wavelets. (Optional)
In parallel, and optionally, perform a short-time Fourier transform (STFT) or a wavelet transform for local context:
@@
    For STFT: Partition the input $\mathbf{X}$ into overlapping windows of length $w$. Let each window be denoted by $\mathbf{X}^{(m)}\in\mathbb{R}^{w\times d}$, where $m$ indexes the window.
@@
  4. 4.
    Alternatively (Wavelets): Decompose each local segment using a chosen wavelet family (e.g. Daubechies) to capture time-varying frequency information at multiple scales.
These local spectra or wavelet coefficients emphasize short-range patterns that might be blurred in a purely global transform. If the application does not require high-resolution local detail, this step can be omitted.
3
Merging Global and Local Representations.
To maximize information flow from both branches, we propose an isometric fusion approach in the frequency (or transform) domain. First, concatenate the global and local representations (spectra or wavelet coefficients) along their feature dimension: [^1] If multiple local windows exist, they may be pooled or aggregated (e.g. average or max) before concatenation, or one can apply a deeper fusion network to combine them.
@@
where $\mathbf{F}_{\mathrm{local}}$ can be a representative fusion of all local windows (or wavelet scales). [^2] If retaining every window or wavelet scale separately, this dimension becomes $n\times(d+\sum_{m}d_{\mathrm{local}}^{(m)})$.
@@
Gating Instead of Concatenation (When Resources Are Limited).
If memory or computational overhead from concatenation is a concern, one may gate the local and global transforms instead. For example, one can learn a scalar or vector gate $\sigma(\mathbf{c})$ (where $\mathbf{c}$ is a global context) that modulates:
$\hat{\mathbf{F}}\;=\;\sigma(\mathbf{c})\,\odot\,\mathbf{F}_{\mathrm{global}}\;+\;\bigl{(}1-\sigma(\mathbf{c})\bigr{)}\,\odot\,\mathbf{F}_{\mathrm{local}}.$
In this case, the fusion dimension is unchanged and can be more lightweight, albeit at the cost of discarding some of the richer information that the concatenation-based approach retains.
4
Adaptive Spectral Filtering.
@@
Here, $\mathbf{F}$ includes both local and global frequency (or wavelet) information from the isometric fusion step (or the gated mixture).
5
Nonlinear Activation (modReLU) and Inverse FFT.
@@
Finally, if using FFT-based local features, we map back to the token domain via the inverse FFT:
$\mathbf{Y}\;=\;\operatorname{IFFT}\bigl{(}\tilde{\mathbf{F}}\bigr{)}\;\in\;\mathbb{R}^{n\times d}.$
For wavelet-based features, one may similarly invert them if a full reconstruction in the token domain is desired. The resulting output is a globally mixed representation that also incorporates local spectral (or wavelet) features (if local transforms were used).
@@ 3.3 Efficient Mixing
Using the FFT to encode global interactions has several advantages. It performs global mixing in $\mathcal{O}(n\log n)$ time, contrasting with the $\mathcal{O}(n^{2})$ cost of self-attention. Meanwhile, local STFT or wavelets capture short-range structure that may be lost in a purely global transform. By isometrically merging (or gating) local and global frequency components, we fuse high-resolution local context with a global view, all in roughly $\mathcal{O}(n\log n)$ time.
@@ 3.6 Proof of Computational Complexity
Local Windowing (STFT) or Wavelets. (Optional)
If each STFT window has length $w$ and we assume an overlap of $\alpha w$ tokens, the number of windows is approximately $\frac{n}{(1-\alpha)w}$. Each windowed FFT thus costs $\mathcal{O}(w\log w)$ for each of the $d$ channels. The total cost is:
$\mathcal{O}\!\Bigl{(}\tfrac{n}{(1-\alpha)w}\;d\;w\log w\Bigr{)}.$
For wavelets, the complexity is usually $\mathcal{O}(n)$ or $\mathcal{O}(n\log n)$ depending on the chosen algorithm and wavelet family, often comparable to STFT approaches. These steps are optional and can be skipped if local structure is not critical for the task.
@@
Figure 1: Overall Process of FFTNet with Local Windowing (STFT or wavelets, optional) and Global FFT, followed by isometric fusion (or gating) in the frequency/transform domain.
@@ 3.7 Comparison with Convolution-Based Methods
Although a sufficiently large convolution kernel can span the entire sequence, it generally lacks the adaptability of self-attention unless the kernel itself is dynamically generated. Our approach bridges this gap by using a data-dependent, learnable filter in the frequency domain, guided by global context and applied through modReLU. Because frequency-domain multiplication corresponds to convolution in the token domain, the model can realize global, context-aware mixing more flexibly than standard convolutional methods. Local windowing or wavelet transforms further refine the representation without resorting to large kernel sizes.
@@ 3.8 Comparison with FNet
FNet Lee-Thorp et al. (2022) similarly uses the DFT for token mixing but employs no learnable parameters in the mixing step. In contrast, our method includes a context-dependent, adaptive filter and a nonlinear activation that both enhance expressivity and allow emphasis on specific frequency components. Further, we incorporate a local STFT or wavelet pathway to capture fine-grained interactions. Despite these adaptive steps, the overall complexity remains $\mathcal{O}(n\log n)$, dominated by the FFT itself.
@@ 3.8.1 Mitigating Complexity in the Adaptive MLP
For very long sequences, generating frequency-bin parameters can still be demanding. Possible solutions include 1 subsampling or pooling the frequency representation before the MLP, 2 sharing parameters among adjacent frequency bins, and 3 applying low-rank factorization to reduce dimensionality. These techniques help maintain a reasonable balance between the expressive power of the adaptive filter and computational cost.
Conclusion. Adaptive spectral filtering integrates frequency-domain global mixing, data-driven filtering, and nonlinear activation into a unified framework. By adding an optional local transform path (STFT or wavelets) and merging (or gating) it with the global FFT via an isometric fusion, the model captures both short-range and long-range interactions with $\mathcal{O}(n\log n)$ efficiency. Parseval’s theorem and the orthogonal properties of the DFT ensure stable transformations, enabling many benefits of self-attention at a reduced computational cost. Full implementation details are available at: https://github.com/jacobfa/fft.
@@ 4 Experiments
We present a comprehensive evaluation of our proposed approach, FFTNet, comparing it against both FNet and traditional Transformers based on self-attention. Our experiments cover the Long Range Arena (LRA) benchmark Tay et al. (2021) and ImageNet classification. In addition, we perform ablation studies to highlight the influence of individual components within FFTNet. We compare against the “standard” variants of the ViT Dosovitskiy et al. (2021) and transformer Vaswani et al. (2017), so as to form an “apples-to-apples” comparison.
@@ 4.2 Long Range Arena (LRA) Benchmark
We evaluate on six LRA tasks: ListOps, Text, Retrieval, Image, Pathfinder, and Path-X. Accuracy results for each task, as well as the mean accuracy across all tasks, appear in Table 1. Here, the standard Transformer serves as the baseline for computing improvements/drops. “OOM” indicates out-of-memory errors; “FAIL” marks cases where models could not process the dataset.
[…]
Table 1: Accuracy (%) on the Long Range Arena (LRA) benchmark tasks. Arrows indicate the performance difference relative to the standard Transformer.
As shown, even the FFTNet (No-Windowing) version outperforms both baselines (Transformer and FNet) on most tasks. Adding local windowing (STFT) further improves the accuracy slightly, achieving the highest average score.
@@ 4.3 ImageNet Classification
We further assess our FFTNetViT variants on ImageNet, comparing them against standard ViT baselines. Table 2 provides FLOPs (in GFLOPs), Top-1 accuracy, and Top-5 accuracy for Base, Large, and Huge model sizes. Although FFTNetViT typically uses fewer FLOPs, the No-Windowing variant is slightly worse than the standard ViT. By contrast, adding local windowing (STFT) brings a small but consistent accuracy boost, surpassing the ViT baseline.
[…]
Table 2: Comparison on ImageNet for Base, Large, and Huge variants. FLOPs are in GFLOPs. “No-Windowing” vs. “With Windowing” are two configurations of FFTNetViT.
@@
Throughput and Speedup Analysis (No-Windowing). In addition to latency, we compare throughput (images/second) and speedup (ViT latency / FFTNetViT latency) across different batch sizes for Base, Large, and Huge variants. Figure 3 displays these two plots side by side, illustrating:
1 Throughput: FFTNetViT variants consistently process more images per second than standard ViT, and the gap widens at larger batch sizes. 2 Speedup: The ratio of ViT latency to FFTNetViT latency is generally above 1, confirming that FFT-based variants are faster. This advantage becomes more pronounced as the batch size increases.
@@ 4.4 Ablation Studies
Ablation studies on ImageNet (Base variant, No-Windowing) were conducted to isolate key FFTNet components. Four configurations are examined. The first is the full FFTNet model (baseline in Table 3), which includes spectral gating, the adaptive module, and FFT-based mixing. The other rows each omit or replace one key component, showing how it affects Top-1 accuracy.
Variant Top-1 Acc (%) Observation FFTNet (full) 79.2 Baseline: spectral gating, adaptive module, FFT-based filtering Without spectral gating 77.8$\downarrow$1.4 Omitting gating causes a moderate performance decline. Without adaptive module 77.2$\downarrow$2.0 Disabling the adaptive module degrades accuracy further. Replace FFT with convolution 75.4$\downarrow$3.8 Using a standard convolution causes the largest drop.
Table 3: Ablation experiments on ImageNet (Base variant, No-Windowing). The FFTNet (full) row is the baseline. Arrows show accuracy drops relative to it.
All components contribute to performance, with substituting the FFT layer for a convolution causing the largest accuracy drop.